If a fair coin is successively flipped, find the probability that a head first appears on the fifth trial.


1
Expert's answer
2021-10-26T14:50:58-0400

In this case we have a geometric distribution


"g(x, p)=pq^{x-1}"

Gicen "p=\dfrac{1}{2}=q, x=5"


"P(X=5)=\dfrac{1}{2}(\dfrac{1}{2})^{5-1}=\dfrac{1}{32}"

The probability that a head first appears on the fifth trial is "\dfrac{1}{32}."